# Meta Pruning via Graph Metanetworks : A Meta Learning Framework for Network Pruning

### v1

@@ Abstract
Network pruning, aimed at reducing network size while preserving accuracy, has attracted significant research interest. Numerous pruning techniques have been proposed over time. They are becoming increasingly effective, but more complex and harder to interpret as well. Given the inherent complexity of neural networks, we argue that manually designing pruning criteria has reached a bottleneck. To address this, we propose a novel approach in which we use a neural network to prune neural networks. More specifically, we introduce the newly developed idea of metanetwork from meta-learning into pruning. A metanetwork is a network that takes another network as input and produces a modified network as output. In this paper, we first establish a bijective mapping between neural networks and graphs, and then employ a graph neural network as our metanetwork. We train a metanetwork that learns the pruning strategy automatically which can transform a network that is hard to prune into another network that is much easier to prune. Once the metanetwork is trained, our pruning needs nothing more than a feedforward through the metanetwork and the standard finetuning to prune at state-of-the-art. Our method achieved outstanding results on many popular and representative pruning tasks (including ResNet56 on CIFAR10, VGG19 on CIFAR100, ResNet50 on ImageNet). Our code is available at https://github.com/Yewei-Liu/MetaPruning.
@@ 1 Introduction
Correspondence to Muhan Zhang
With the rapid advancement of deep learning [33, 53], neural networks have become increasingly powerful. However, this improved performance often comes with a significant increase in the number of parameters and computational cost (FLOPs). As a result, there is growing interest in methods to simplify these networks while preserving their performance. Pruning, which involves selectively removing certain parts of a neural network, has proven to be an effective approach [2, 17, 51].
A key idea of a large number of previous pruning works is to remove the unimportant components of a neural network. To achieve this goal, various criteria that measure the importance of components of a neural network have been invented. Some criteria are based on norm or magnitude scores [15, 36, 55, 18]. Some are based on how much the performance changes when certain parts of the network are removed or preserved [68, 46, 10, 66]. Some others defined more complex criteria such as saliency and sensitivity [32, 34, 72].
Once a pruning criterion is established, we can assess whether a model is easy or difficult to prune based on that criterion. For instance, using a simple $l_{2}$ norm criterion, a model whose parameters exhibit either very large or very small $l_{2}$ norms is clearly easier to prune compared to one with uniformly distributed $l_{2}$ norms. From this viewpoint, we can classify models as more or less suitable for pruning. This concept forms the foundation of our work.
Previously, researchers have generally approached pruning improvement in two ways: either by refining the pruning criteria or by transforming the model from “bad” to “good” according to a given criterion. The first approach is more straightforward and has been extensively studied in prior work. In contrast, the second approach has received comparatively less attention. However, several studies adopting this latter perspective have achieved promising results. Notably, sparsity regularization methods [61, 9, 60] exemplify this approach by encouraging models to become sparser, thereby facilitating more effective pruning.
Our work follows the second approach and introduces a novel method inspired by sparsity regularization. We focus on this approach because, despite the many pruning criteria proposed, most lack solid theoretical foundations and have become overly complex. Therefore, in this work, we adopt a simple variant of the $l_{2}$ norm—the group norm from Fang et al. [9]—as our pruning criterion. Additionally, we bring the concept of metanetworks from meta-learning into pruning. A metanetwork is a network that takes another neural network as input and produces a modefied neural network as output. In our method, the metanetwork transforms a “bad” network into a “good” one, making it more amenable to pruning. Once trained, our metanetwork-based approach requires only a feedforward pass through the metanetwork and some standard finetuning to achieve state-of-the-art pruning results.
@@
  • Introducing the idea of metanetwork in meta-learning into pruning for the first time.
  • Proposing a novel pruning method that is simple and powerful.
  • Achieving state-of-the-art performance on several popular pruning tasks.
It is worth noting that this paper focuses solely on structural pruning, which does not rely on specialized AI accelerators or software to reduce memory usage and computational costs. This makes it more versatile and broadly applicable. However, all our methods can also be applied directly to unstructured pruning.
@@ 2.1 Metanetworks
Using neural networks to process other neural networks has emerged as an intriguing research direction. Early studies demonstrated that neural networks can extract useful information directly from the weights of other networks [58]. We define a metanetwork as a neural network that takes another neural network as input and outputs either information about it or a modified network. Initial metanetworks were simple multilayer perceptrons (MLPs), while more recent designs incorporate stronger inductive biases by preserving symmetries [13] inherent in the input networks. Broadly, metanetwork architectures can be categorized into two perspectives. The weight space view applies specially designed MLPs directly on the model’s weights [71, 48, 73, 57, 74, 75, 56]. In contrast, the graph view transforms the input network into a graph and applies graph neural networks (GNNs) to it [40, 30, 27]. In this work, we adopt the graph view to build our metanetwork, leveraging the natural correspondence between graph nodes and neurons in a network layer, as well as the symmetries they exhibit [47]. Given the maturity and effectiveness of GNNs, the graph metanetwork approach offers both elegance and practicality.
@@ 2.2 Graph Neural Networks
Graph neural networks (GNNs) are a class of neural networks specifically designed to operate on graph-structured data, where graphs consist of nodes and edges with associated features [63, 52, 29]. GNNs have gained significant attention in recent years, with frameworks such as the message passing neural network (MPNN) [12] proving to be highly effective. In this paper, we employ the message passing framework, specifically using Principal Neighborhood Aggregation (PNA) [3] as the backbone architecture for our metanetwork.
2.3 Sparsity Regularization-Based Pruning
Networks that are inherently sparse tend to be easier to prune effectively. Prior research has explored various regularization techniques to encourage sparsity in neural networks to facilitate pruning. This idea of sparsity regularization has been widely adopted across many works [61, 14, 25, 44, 22, 43, 64, 65, 9, 60]. Our method inherits this principle to some extent, as our metanetwork can be viewed as transforming the original network into a sparser version that is more amenable to pruning.
@@ 2.4 Learning to Prune & Meta Pruning
Given the complexity of pruning, several previous works have leveraged neural networks [5, 22, 62, 1], reinforcement learning [50, 70, 23], and other techniques to automatically learn pruning strategies. Viewing pruning as a learning problem naturally leads to the idea of meta-learning, which learns “how to learn” [24]. Some prior works have specifically applied meta-learning to pruning [45, 37]. While our work also adopts a meta-learning framework, it differs fundamentally from these previous approaches by using a metanetwork.
3 Method
Our approach centers on training a metanetwork that learns the pruning strategy automatically and can transform a neural network that is difficult to prune into one that is easier to prune, thereby improving pruning effectiveness and efficiency.
We begin by describing how our metanetwork transforms one network into another, including the conversion between network and graph, and the design of our metanetwork(a graph neural network). Next, we introduce the group $l_{2}$ norm criterion for meta-training and pruning. Then we detail the meta-training pipeline used to train the metanetwork. Finally, we present the complete pruning process.

### v2

# Meta Pruning via Graph Metanetworks : A Universal Meta Learning Framework for Network Pruning

@@ Abstract
We propose an entirely new meta-learning framework for network pruning. It is a general framework that can be theoretically applied to almost all types of networks with all kinds of pruning and has great generality and transferability. Experiments have shown that it can achieve outstanding results on many popular and representative pruning tasks (including both CNNs and Transformers). Unlike all prior works that either rely on fixed, hand-crafted criteria to prune in a coarse manner, or employ learning to prune ways that require special training during each pruning and lack generality. Our framework can learn complex pruning rules automatically via a neural network (metanetwork) and has great generality that can prune without any special training. More specifically, we introduce the newly developed idea of metanetwork from meta-learning into pruning. A metanetwork is a network that takes another network as input and produces a modified network as output. In this paper, we first establish a bijective mapping between neural networks and graphs, and then employ a graph neural network as our metanetwork. We train a metanetwork that learns the pruning strategy automatically and can transform a network that is hard to prune into another network that is much easier to prune. Once the metanetwork is trained, our pruning needs nothing more than a feedforward through the metanetwork and some standard finetuning to prune at state-of-the-art. Our code is available at https://github.com/Yewei-Liu/MetaPruning.
Figure 1: For a pruning criterion, we use a metanetwork to change a hard to prune network into another easy to prune network for better pruning.
@@ 1 Introduction
With the rapid advancement of deep learning (LeCun et al., 2015; Schmidhuber, 2015), neural networks have become increasingly powerful. However, this improved performance often comes with a significant increase in the number of parameters and computational cost (FLOPs). As a result, there is growing interest in methods to simplify these networks while preserving their performance. Pruning, which involves selectively removing certain parts of a neural network, has proven to be an effective approach (Cheng et al., 2024; He & Xiao, 2024; Reed, 1993).
A key idea of a large number of previous pruning works is to remove the unimportant components of a neural network. To achieve this goal, various criteria that measure the importance of components of a neural network have been invented. Some criteria are based on norm or magnitude scores (Han et al., 2016; Li et al., 2017b; Sun et al., 2024; He et al., 2018a). Some are based on how much the performance changes when certain parts of the network are removed or preserved (You et al., 2019b; Ma et al., 2023; Frankle & Carbin, 2019; Ye et al., 2020). Some others defined more complex criteria such as saliency and sensitivity (LeCun et al., 1989; Lee et al., 2019; Zhao et al., 2019).
Once a pruning criterion is established, we can assess whether a model is easy or hard to prune based on that criterion. Technically, we consider a model easy to prune if we can prune a relatively large part of it with only little loss in accuracy. Otherwise, we consider it hard to prune.
Prior researchers have generally approached pruning improvement in two ways: either by refining the pruning criteria or by transforming the model to be easier to prune according to a given criterion. The first approach is more straightforward and has been extensively studied in prior work. In contrast, the second approach has received comparatively less attention. However, several studies adopting this latter perspective have achieved promising results. Notably, sparsity regularization methods (Wen et al., 2016; Fang et al., 2023; Wang et al., 2021) exemplify this approach by encouraging models to become sparser, thereby facilitating more effective pruning.
Our framework inherits the idea of the second way. A metanetwork is a network that takes another network as input and produces a modefied network as output. For any given pruning criterion, our framework aims at training a metanetwork that can transform a hard to prune network into another easier to prune network like figure 1.
@@
  1. (1)
    Introduce the idea of metanetwork from meta-learning into pruning for the first time.
  2. (2)
    Propose an entirely new universal pruning framework that is theoretically applicable to almost all types of networks with all kinds of pruning.
  3. (3)
    Present concrete implementations for our theoretical framework and achieve state-of-the-art performance on various practical pruning tasks.
  4. (4)
    Conduct further research on the flexibility and generality of our framework.
Our method is entirely new and fundamentally different from all prior works, see Appendix A for a comparison with prior works to better understand our novel contributions and advantages.
@@ 2.1 Metanetworks
Using neural networks to process other neural networks has emerged as an intriguing research direction. Early studies demonstrated that neural networks can extract useful information directly from the weights of other networks (Unterthiner et al., 2020). We define a metanetwork as a neural network that takes another neural network as input and outputs either information about it or a modified network. Initial metanetworks were simple multilayer perceptrons (MLPs), while more recent designs incorporate stronger inductive biases by preserving symmetries (Godfrey et al., 2022) inherent in the input networks. Broadly, metanetwork architectures can be categorized into two perspectives. The weight space view applies specially designed MLPs directly on the model’s weights (Zaheer et al., 2017; Navon et al., 2023; Zhou et al., 2023a; Tran et al., 2024b; Zhou et al., 2023b; 2024; Tran et al., 2024a). And the graph view transforms the input network into a graph and applies graph neural networks (GNNs) to it (Lim et al., 2024; Kofinas et al., 2024; Kalogeropoulos et al., 2024). In this work, we adopt the graph view to build our metanetwork, leveraging the natural correspondence between graph nodes and neurons in a network layer, as well as the symmetries they exhibit (Maron et al., 2019). Given the maturity and effectiveness of GNNs, the graph metanetwork approach offers both elegance and practicality.
@@ 2.2 Graph Neural Networks
Graph neural networks (GNNs) are a class of neural networks specifically designed to operate on graph-structured data, where graphs consist of nodes and edges with associated features (Wu et al., 2021; Scarselli et al., 2009; Kipf & Welling, 2017). GNNs have gained significant attention in recent years, with frameworks such as the message passing neural network (MPNN) (Gilmer et al., 2017) proving to be highly effective. In this paper, we employ the message passing framework, specifically using Principal Neighborhood Aggregation (PNA) (Corso et al., 2020) as the backbone architecture for our metanetwork.
2.3 Sparsity Regularization Based Pruning
Networks that are inherently sparse tend to be easier to prune effectively. Prior research has explored various regularization techniques to encourage sparsity in neural networks to facilitate pruning. This idea of sparsity regularization has been widely adopted across many works (Wen et al., 2016; Gordon et al., 2018; Huang & Wang, 2018; Lin et al., 2020c; He et al., 2017; Lin et al., 2019; Xia et al., 2022; 2024; Fang et al., 2023; Wang et al., 2021). Our method inherits this principle to some extent, as our metanetwork can be viewed as transforming the original network into a sparser version for better pruning.
@@ 2.4 Learning to Prune & Meta Pruning
Given the complexity of pruning, several previous works have leveraged neural networks (Dery et al., 2024; He et al., 2017; Wu et al., 2024; Chen et al., 2023), reinforcement learning (Rao et al., 2019; Yu et al., 2022; He et al., 2018b), and other techniques to automatically learn pruning strategies. Viewing pruning as a learning problem naturally leads to the idea of meta-learning, which learns “how to learn” (Hospedales et al., 2022). Some prior works have specifically applied meta-learning to pruning (Liu et al., 2019; Li et al., 2020). Our work also draws idea from meta-learning, but it learns in a way entirely different from all prior works.
3 A Universal Meta-Learning Framework
Our framework can be summarized in one single sentence (Figure 1):
For a pruning criterion, we use a metanetwork to change a hard to prune network into another easy to prune network for better pruning.
Pruning criterion: A pruning criterion measures the importance of specific components of a neural network. Typical examples include the $\ell_{1}$ norm and $\ell_{2}$ norm of weight vectors. During pruning, we compute an importance score for each prunable component according to the criterion, and remove components in ascending order of their scores.
Easy or hard to prune: For a fixed pruning criterion, if we gradually prune the network and the accuracy drops quickly, we refer to it as hard to prune. Conversely, if the accuracy decreases slowly with progressive pruning, we call it easy to prune.
Metanetwork: A metanetwork is a special type of neural network whose inputs and outputs are both neural networks. In contrast, standard neural networks typically take structured data such as images (CNNs), sentences (Transformers), or graphs (GNNs) as input. A metanetwork, however, takes an existing network as input and outputs another network, hence the term meta—indicating that it operates on neural networks themselves.
4 A Specific Implementation Based on Graph Metanetworks
The proposed framework is conceptually simple and straightforward. However, its practical implementation requires addressing several technical challenges:
  1. (1)
    How to build the metanetwork ?
  2. (2)
    How to train the metanetwork ?
  3. (3)
    How to design the pruning criterion ?
We present our implementations step by step, dedicating one subsection to each challenge. The final subsection offers an overall summary of our method and a deeper understanding of it from a more holistic perspective.
4.1 Metanetwork Design
Drawing ideas from recent researches (section 2.1), we use a graph neural network (GNN) as our metanetwork. Please make sure you are familiar with the basic concepts of GNN before continuing. Our design does not cover any complex GNN theory, you only need to know what is a graph and the basic message passing algorithm.
Figure 2: Conversion between networks and graphs: We visualize a two layer toy CNN as example. Our graph contains 6 nodes corresponding with 6 neurons in the network. For neurons connected via convolutional channel like (1,3) and connected by residual connection like (1,5), we add an edge between their corresponding nodes. Then we generate node and edge features as shown in the figure. For some special cases, we use default values to replace non-existent values. For example, for node 3 who doesn’t have a residual connection, we set the last 4 number of its node feature as [1, 0, 0, 1]. This means we treat it as with previous residual batchnorm of weight 1, bias 0, running mean 0, running var 1, performing the same as with no batchnorm.
Conversion between networks and graphs: To apply a metanetwork (GNN) on networks, we first need to establish a conversion between networks and graphs. Prior works such as Kofinas et al. (2024); Lim et al. (2024) have managed to change networks of almost all architectures (CNN, Transformer, RNN, etc.) into graphs.
Here we show how we establish conversion between a ResNet (He et al., 2016) and a graph in our experiments as an example. This inherits the idea of Kofinas et al. (2024); Lim et al. (2024) but is quite different in implementation details. To convert a network into a graph, we establish the correspondence between the components of a neural network and elements of a graph as follows:
(1) Node: Each neuron in fully connected layers or channel in convolutional layers in the network is represented as a node in the graph. (2) Edge: An edge exists between two nodes if there is a direct connection between them in the original network. This includes fully connected layers, convolutional layers, and residual connections. (3) Node Features: Node features comprise parameters associated with neurons, here we use the weight, bias, running mean, running var of batchnorm (Ioffe & Szegedy, 2015), including both batchnorms from previous adjacent layers and previous skip connection layers (if exist). (4) Edge Features: Edge features encode the weights of connections between nodes. We treat linear connections and residual connections as special cases of convolutional connections with a kernel size of 1. For a $k\times k$ convolution between two channels, we flatten it into a $k^{2}$-dim edge feature vector.
Since graph neural networks here require all edge features to have the same dimension, we adopt one of two strategies to standardize them: (1) Pad all convolutional kernels to the same size (e.g., the maximum kernel size in the network) (2) Flatten convolutional kernels of varying sizes, then apply learned linear transformations to project them into the same size.
During conversion, we simply ignored other components such as pooling layers. Although we have methods to explicitly convert all of them, our experiments show that omitting them already yields outstanding results. Therefore, it is acceptable to ignore them in practice. We also didn’t use any other techniques like positional embedding.
Metanetwork architecture: After establishing the conversion between networks and graphs, we can convert the input network into a graph, use metanetwork(GNN) to transform the graph, and finally convert the output graph back into a new network. Our metanetwork must be powerful enough to learn the transforming rules for better pruning. We build our metanetwork based on the message passing framework (Gilmer et al., 2017) and PNA (Corso et al., 2020) architecture. A brief overview of our architecture is provided below, the full description is available in Appendix B.1.
$v\leftarrow\text{MLP}_{NodeEnc}(v_{in}),$ (1)
$e\leftarrow\text{MLP}_{EdgeEnc}(e_{in}).$ (2)
Then they will pass through several message passing layers. In each layer, we update node and edge features using information from adjacent nodes and edges.
$v_{i}$ $\leftarrow f_{v}(v_{i},v_{j}((i,j)\in E),e_{ij}((i,j)\in E))$ (3)
$e_{ij}$ $\leftarrow f_{e}(v_{i},v_{j},e_{ij})$ (4)
Here $f_{v}$ and $f_{e}$ are fixed calculating processes.
$v_{pred}\leftarrow\text{MLP}_{NodeDec}(v),$ (5)
$e_{pred}\leftarrow\text{MLP}_{EdgeDec}(e).$ (6)
We multiply predictions by a residual coefficient and add them to the original inputs as final outputs:
$v_{out}=\alpha\cdot v_{pred}+v_{in},$ (7)
$e_{out}=\beta\cdot e_{pred}+e_{in},$ (8)
4.2 Meta-Training
Meta-training is the process we train our metanetwork (Figure 3).
Figure 3: Meta-Training Pipeline: During each iteration, the origin network is converted into the origin graph, fed through the metanetwork to get a new graph and finally converted back into a new network. We calculate accuracy loss and sparsitly loss on the new network, then backpropagated the gradients to update the metanetwork.
Data preparation: Meta-training data is comprised of two parts. The first is a set of neural networks to prune, which serve as the origin network in meta-training. We call them data models. The number of data models can be arbitrary small numbers like 1, 2, or 8, because our implementations aren’t sensitive to it (See more explanations in Section 4.4).The second is the traditional training datasets, CIFAR-10 for example, used to calculate the accuracy loss.
One training iteration: We select one model from data models as the origin network. It is converted into its graph representation, passed through the metanetwork to get a new graph and converted back into a new network. We calculate accuracy loss and sparsity loss on this new network, and backpropagated the gradients to update the metanetwork.
Two types of losses are used during meta-training.(1) Accuracy Loss: We feed the training data (e.g., Train set of CIFAR-10. Note that test set is not used here) into the new network and compute the cross-entropy loss based on the output predictions, ensuring the metanetwork doesn’t excessively disturbing the effective parts of the origin network. (2) Sparsity Loss: We calculate a regularization term on the new network to encourage it to be sparse, making it easier to prune. It is related to the design of pruning criterion (section 4.3)
4.3 Pruning Criterion
A valid pruning criterion consists of two components:(1) An importance score function: During pruning, we calculate the scores of all prunable components of a network, and prune them in ascending order of their scores.(2) A sparsity loss: It is a kind of loss that is calculated on the parameters of the network. Optimizing the network using this loss can make it easier to prune.
We draw ideas from prior sparsity regularization based pruning methods(section 2.3), and design our pruning criterion based on them (especially Fang et al. (2023)). We will give a brief introduction below. For a rigorous mathematical definition and more details, please refer to appendix B.2.
We only do structural pruning in all our experiments. A variant of the classical $\ell_{2}$ norm, the group norm is used as our importance score function. It inherits the idea of using $\ell_{2}$ norm to calculate a score, but calculate it on a pruning group level. Since structural pruning is performed on groups, traditional $l_{2}$ norm that treats nodes and edges in isolation is no longer viable. Group $l_{2}$ norm that takes a whole group into consideration has shown to be more effective. We calculate the sparisty loss by multiply our group norm scores with some coefficients.
The choice of pruning criteria is highly flexible (see in our later experiments, section 6.1), which further demonstrates the generality of our framework. Although we conduct only structural pruning in all our experiments, all our methods can be directly applied to unstructured pruning as well.
4.4 A Holistic Perspective
Figure 4: Feed forward through metanetwork: when a network is fed forward through the metanetwork, every parameter in it gathers information from neighbour parameters and architectures.
Intuitively, when a network is fed forward through the metanetwork, every parameter in it gathers information from neighbour parameters and architectures (Figure 4). The metanetwork automatically learns how to process the information from neighbours and change the weights of the network based on that for better pruning. And this has several important properties:
Generality: Our metanetwork is a really small GNN compared to the origin network, which means each parameter can only see information from a limited local region. Due to the nature of GNN, parameters at any position in the network—whether at the beginning, middle, or end—must process information from neighbours in the same way. As a result, the metanetwork must learn a general strategy to adjust parameters based on neighbour information for better pruning. This explains why, in all our experiments, the pruning models and meta-training data models are completely different, but the performance is as good as the same. This also explains why our implementations are not sensitive to the number of data models. Even a single data model can be viewed as a collection of numerous data points sufficient to train the metanetwork. So there is no overfitting at all, and whether we are using 1, 2 or 8 data models during meta-training, the results remain the same.
Natural transferability: Just as a GNN can be applied to graphs of varying sizes without modification, our metanetwork can be directly applied to networks of different scales. This property gives it natural transferability, as demonstrated in our experiments (section 6.2) where it successfully transfers across related datasets and network architectures.
Universally applicability: Theoretically, our implementations can be applied to almost all types of networks with all kinds of pruning. Almost all types of networks can be converted into graphs (Kofinas et al., 2024; Lim et al., 2024), and the designs of the GNN and the pruning criteria are also broadly compatible. This makes our framework a universally applicable solution for network pruning.
5 Experiments on Classical CNN Pruning Tasks
5.1 Preliminaries
$\underbrace{\text{Initial Pruning (Finetuning)}}_{\text{Optional}}\rightarrow\underbrace{\text{Metanetwork (Finetuning)}\rightarrow\text{Pruning (Finetuning)}}_{\text{Necessary}}$ (9)
Definitions of the terms Speed Up and Acc vs. Speed Up Curve are in Appendix C.1. The pruning pipeline is described as equation 9, and full details are provided in Appendix C.2.
Unlike prior learning to prune methods that require special training during each pruning, our pruning pipeline only requires a feed forward through metanetwork and standard finetuning, which is simple, general and effective. Once we have a metanetwork, we can prune as many networks as we want. All pruning models are completely different from meta-training data models. This demonstrates our metanetwork naturally has great transferability and no prior work has done something like this before. See Appendix A for more comparisons between our work and prior ones.
5.2 Classical CNN Pruning Tasks
We carry out our experiments on three most classical, popular, and representative image recognition tasks, including pruning ResNet56 (He et al., 2016) on CIFAR10 (Krizhevsky & Hinton, 2009), VGG19 (Simonyan & Zisserman, 2015) on CIFAR100 (Krizhevsky & Hinton, 2009) and ResNet50 (He et al., 2016) on ImageNet (Deng et al., 2009). See appendix D.1 for more general setups.
Our method achieves outstanding results on all 3 tasks and is better than almost all prior works (Table 1). See full results compared with prior works in Table 6(ResNet56 on CIFAR10), Table 9(VGG19 on CIFAR100), Table 12(ResNet50 on ImageNet). See Appendix D.2 D.3 D.4 for implementation details on each tasks.
Table 1: Results for 3 classical CNN pruning tasks including pruning (1) ResNet56 on CIFAR10, (2) VGG19 on CIFAR100 (3) ResNet50 on ImageNet. For full results compared with prior works, see (1) Table 6, (2) Table 9, (3) Table 12.
Task Base Top-1(Top-5) Pruned Top-1($\Delta$) Pruned Top-5($\Delta$) Pruned FLOPs
(1) 93.51% 93.64%(+0.13%) 65.6%
(2) 73.65% 69.75%(-3.90%) 88.83%
(3) 76.14%(93.11%) 76.13%(-0.01%) 92.78%(-0.33%) 57.2%
5.3 Ablation Study
How metanetwork works: We visualized the "Acc vs. Speed Up" curve of both origin network and network after feedforward through metanetwork and finetuning (Figure 5). They show that as the pruning speed up increases, the accuracy drops at a significantly slower rate after applying the metanetwork and finetuning.
(a) ResNet56 on CIFAR10
(b) VGG19 on CIFAR100
(c) ResNet50 on ImageNet
(d) ViT on ImageNet
Figure 5: Metanetwork changes hard to prune network into easy to prune network.
Trade-off between accuracy and speed up: During meta-training: as the number of training epochs increases, the metanetwork’s ability to make the network easier to prune becomes stronger. However, its ability to maintain the accuracy becomes weaker. As shown in Figure 6(a), with more training epochs, the "Acc VS. Speed Up" curve shifts downward, and the flat portion of the curve becomes longer. This characteristic allows us to adaptively meet different pruning requirements. If a higher level of pruning is desired and a moderate drop in accuracy is acceptable, a metanetwork trained for more epochs would be preferable. Conversely, if preserving accuracy is more important, a metanetwork trained for fewer epochs would be a better choice (Figure 6(b)).
(a) Acc VS. Speed Up curves of metanetworks from different meta-training epochs
(b) Feed network through metanetworks from different meta-training epochs, finetuning, then prune it progressively to find the maximum speed up that can maintain the accuracy above a certain threshold.
Figure 6: Trends in meta-training. (ResNet56 on CIFAR10 as example)
Finetuning after Metanetwork. See appendix E.1 for how number of finetuning epochs after metanetwork influence the pruning results.
6 Experiments on Transferability and Flexibility
6.1 Flexible Pruning Criterion
Pruning criterion of our method is highly flexible. We build a series of reasonable pruning criterion (Appendix B.2). At first, we directly use MEAN REDUCE, MAX NORMALIZE, $\alpha=4$ as default, because this is the same as group norm in Fang et al. (2023). Later, we keep everything else the same, and try many different criteria, and find they all works well (Table 2). This demonstrates that our framework is robust and has many flexibilities.
Table 2: Flexible pruning criterion: Pruning ResNet56 on CIFAR10, all experiments use the same origin network wtih Test Acc 93.51%, all criteria use MEAN REDUCE. We tried different values of Alpha and different ways of NORMALIZE, they all work well. Even in the Naive situation, where we use no NORMALIZE and no shrinkage strength (Alpha is 0), the results remain robust.
NORMALIZE alpha Pruned Acc Pruned FLOPs Speed Up
Default MAX 4 93.64% 65.64% 2.91
Alpha MAX 0 93.37% 65.75% 2.92
MAX 2 92.87% 68.15% 3.14
MAX 6 93.36% 65.75% 2.92
NORMALIZE MEAN 4 93.42% 65.52% 2.90
NONE 4 93.08% 65.64% 2.91
Naive NONE 0 93.04% 65.75% 2.92
6.2 Transfer between Datasets and Architectures
Section 4.4 mentioned that our metanetwork has natural transferability. Following experiments demonstrate that our metanetwork can transfer between similar datasets and network architectures.
Transfer between datasets: 3 datasets, CIFAR10 (Krizhevsky & Hinton, 2009), CIFAR100 (Krizhevsky & Hinton, 2009) and SVHN (Goodfellow et al., 2014) are used. We train metanetwork and the to be pruned network on each of them, traverse through every possible combination, and use the metanetwork to prune the network. Results (Table 3) show that out metanetwork can transfer between similar datasets. See Appendix F.1 for full details
Table 3: Transfer between datasets: All networks’ architecture is ResNet56. Columns represent the training datasets for the metanetwork, and rows represent the training datasets for the to be pruned network. “None” indicates using no metanetwork. Results with metanetwork is obviously better than no metanetwork (The only exception is when training datasets for the to be pruned network is SVHN, and we guess this is because the dataset SVHN itself is too easy).